# Enskog kinetic theory for a model of a confined quasi-two-dimensional granular fluid

### v6

@@ I Introduction
$\kappa_{\text{k}}^{*}=\frac{1}{2\nu_{\kappa}^{*}+\Delta_{s}^{*}\Big{(}\frac{\partial\zeta_{0}^{*}}{\partial\Delta^{*}}\Big{)}_{s}}\left\{1+\frac{3}{8}\phi\chi(1+\alpha)^{2}(2\alpha-1)-\frac{\Delta_{\text{s}}^{*}}{\sqrt{2\pi}}\phi\chi\left[\frac{7}{4}+2(1+\alpha)\left(1-\frac{3}{4}\sqrt{2\pi}\Delta_{\text{s}}^{*}\right)-3(1+\alpha)^{2}-\Delta_{\text{s}}^{*2}\right]\right\}$,
@@
$\mu_{\text{k}}^{*}=-\frac{1}{\nu_{\kappa}^{*}}\phi\chi\left(1+\frac{1}{2}\phi\partial_{\phi}\ln\chi\right)\left\{\frac{3}{8}\alpha(1-\alpha^{2})-\frac{\Delta_{\text{s}}^{*}}{\sqrt{2\pi}}\phi\chi\left[2(\Delta_{\text{s}}^{*2}+\alpha^{2})-\frac{1}{2}+\frac{3}{2}\sqrt{2\pi}\alpha\Delta_{\text{s}}^{*}\right]\right\}$,
@@ II.2 Enskog kinetic equation
×$\times(-\widehat{\boldsymbol{\sigma}}\cdot{\bf g}-2\Delta)\alpha^{-2}\chi(\mathbf{r},\mathbf{r}+\boldsymbol{\sigma})f(\mathbf{r},\mathbf{v}_{1}^{\prime\prime},t)$
×$\times f(\mathbf{r}+\boldsymbol{\sigma},\mathbf{v}_{2}^{\prime\prime},t)-\sigma^{d-1}\int\ \text{d}{\bf v}_{2}\int\text{d}\widehat{\boldsymbol{\sigma}}\Theta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})$
×$\times(\widehat{\boldsymbol{\sigma}}\cdot{\bf g})\chi(\mathbf{r},\mathbf{r}+\boldsymbol{\sigma})f(\mathbf{r},\mathbf{v}_{1},t)f(\mathbf{r}+\boldsymbol{\sigma},\mathbf{v}_{2},t).$ (9)
@@
×$\times\chi(\mathbf{r},\mathbf{r}+\boldsymbol{\sigma})f(\mathbf{r},\mathbf{v}_{1},t)f(\mathbf{r}+\boldsymbol{\sigma},\mathbf{v}_{2},t)\left[\psi(\mathbf{v}_{1}^{\prime})-\psi(\mathbf{v}_{1})\right],$
@@ III.1 Zeroth-order approximation
×$\times(-\widehat{\boldsymbol{\sigma}}\cdot{\bf g}-2\Delta)\alpha^{-2}f^{(0)}(\mathbf{v}_{1}^{\prime\prime})f^{(0)}(\mathbf{v}_{2}^{\prime\prime})-\sigma^{d-1}\chi$
×$\times\int\ \text{d}{\bf v}_{2}\int\text{d}\widehat{\boldsymbol{\sigma}}\Theta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})(\widehat{\boldsymbol{\sigma}}\cdot{\bf g})f^{(0)}(\mathbf{v}_{1})f^{(0)}(\mathbf{v}_{2}).$ (29)
@@
$-\zeta^{(0)}T\partial_{T}f^{(0)}=J_{\text{E}}^{(0)}[f^{(0)},f^{(0)}].$ (32)
@@
×$\times\left(B_{1}g^{*}\Delta^{*2}+B_{2}\alpha g^{*2}\Delta^{*}-\frac{1-\alpha^{2}}{4}B_{3}g^{*3}\right),$
@@ III.2 First-order approximation
The quantities $\boldsymbol{\mathcal{A}}$, $\boldsymbol{\mathcal{B}}$, $\mathcal{C}_{ij}$ and $\mathcal{D}$ are the solutions of the following linear integral equations:
@@
${\bf A}\left(\mathbf{V}\right)=$ $-\mathbf{V}T\frac{\partial f^{(0)}}{\partial T}-\frac{p}{\rho}\frac{\partial f^{(0)}}{\partial\mathbf{V}}-\boldsymbol{\mathcal{K}}\left[T\frac{\partial f^{(0)}}{\partial T}\right],$ (47)
${\bf B}\left(\mathbf{V}\right)=$ $-{\bf V}f^{(0)}-\frac{p}{\rho}\left(1+\phi\frac{\partial}{\partial\phi}\ln p^{*}\right)\frac{\partial f^{(0)}}{\partial\mathbf{V}}$
$-\left(1+\frac{1}{2}\phi\frac{\partial}{\partial\phi}\ln\chi\right)\boldsymbol{\mathcal{K}}\left[f^{(0)}\right],$ (48)
@@
$D\left(\mathbf{V}\right)=$ $\frac{1}{d}\frac{\partial}{\partial\mathbf{V}}\cdot\left(\mathbf{V}f^{(0)}\right)+\left(\zeta_{U}+\frac{2}{d}p^{*}\right)T\frac{\partial f^{(0)}}{\partial T}$
$+\frac{1}{d}\mathcal{K}_{i}\left[\frac{\partial f^{(0)}}{\partial V_{i}}\right].$ (50)
@@
×$\times(-\widehat{\boldsymbol{\sigma}}\cdot\mathbf{g}-2\Delta)\widehat{\sigma}_{i}\alpha^{-2}f^{(0)}(\mathbf{v}_{1}^{\prime\prime})X(\mathbf{v}_{2}^{\prime\prime})$
@@
×$\times f^{(0)}(\mathbf{v}_{1})X(\mathbf{v}_{2}).$
@@ IV.1 Pressure tensor
$I_{\eta}$ $=\int\text{d}\mathbf{c}_{1}\int\text{d}\mathbf{c}_{2}\;g^{*-1}g_{x}^{*2}g_{y}^{*2}\varphi_{\text{M}}(\mathbf{c}_{1})\varphi(\mathbf{c}_{2}),$ (57)
@@ IV.2 Heat flux
×$\times\frac{1}{dn}\int\;\text{d}\mathbf{v}\mathbf{S}(\mathbf{V})\cdot\boldsymbol{\mathcal{K}}\left[f^{(0)}\right],$
@@
As mentioned before, the explicit dependence of the (reduced) Navier-Stokes transport coefficients on the solid volume fraction and the coefficient of restitution requires knowledge of the quantities $\zeta^{(0)}$, $a_{2}$, $\nu_{\eta}$, $\nu_{\kappa}$, and $\nu_{\mu}$, the dimensionless integrals $\left\{I_{\eta},I_{\gamma},I_{\eta}^{\prime},I_{\kappa},I_{\kappa}^{\prime},I_{\kappa}^{\prime\prime}\right\}$, and two integrals involving the operator $\boldsymbol{\mathcal{K}}$. All these quantities are given in terms of the zeroth-order scaled distribution $\varphi$ and the solutions $\boldsymbol{\mathcal{A}}$, $\boldsymbol{\mathcal{B}}$, and $\mathcal{C}_{ij}$ to the linear integral equations (65), (IV.2), and (82).
@@
$\varphi(\mathbf{c},\Delta^{*})$$\approx$ $\varphi_{\text{M}}(\mathbf{c})\left\{1+a_{2}(\Delta^{*})\left[\frac{c^{4}}{2}-\frac{d+2}{2}c^{2}\right.\right.$ (84)
@@
$I_{\eta}=\frac{3}{8}\sqrt{\frac{\pi}{2}},\quad I_{\gamma}=\sqrt{\frac{\pi}{2}},\quad{I}_{\eta}^{\prime}=-\sqrt{2\pi}\left(\frac{3}{4}+\alpha\right),$ (88)
@@
Finally, to evaluate the kinetic parts of the transport coefficients, one still needs to know the explicit forms of the collision frequencies $\nu_{\eta}$, $\nu_{\kappa}$ and $\nu_{\mu}$ as well as the integrals involving the operator $\boldsymbol{\mathcal{K}}[X]$. In the case of the collision frequencies, one takes the approximations (115) for $\mathcal{C}_{ij}(\mathbf{V})$ and (121) for $\boldsymbol{\mathcal{A}}$ and $\boldsymbol{\mathcal{B}}$. These integrals have been performed in Ref. BBMG15 for a $d$-dimensional system. In the case $d=2$, one gets
@@
$\int\text{d}{\bf V}{\bf S}({\bf V})\cdot\boldsymbol{\mathcal{K}}\left[\frac{\partial}{\partial\mathbf{V}}\cdot\left(\mathbf{V}f^{(0)}\right)\right]=$ $8\phi\chi\frac{nT^{2}}{m}\bigg{\{}\frac{3}{8}(1+\alpha)^{2}(1-2\alpha)+\frac{\Delta^{*}}{\sqrt{2\pi}}\bigg{[}\frac{7}{4}+2(1+\alpha)\left(1-\frac{3}{4}\sqrt{2\pi}\Delta^{*}\right)$
$-3(1+\alpha)^{2}-\Delta^{*2}\bigg{]}\bigg{\}},$ (93)
$\int\text{d}\mathbf{v}\mathbf{S}(\mathbf{V})\cdot\boldsymbol{\mathcal{K}}\left[f^{(0)}\right]=$ $-8\phi\chi\frac{nT^{2}}{m}\left\{\frac{3}{8}\alpha(1-\alpha^{2})-\frac{\Delta^{*}}{\sqrt{2\pi}}\left[2(\Delta^{*2}+\alpha^{2})-\frac{1}{2}+\frac{3}{2}\sqrt{2\pi}\alpha\Delta^{*}\right]\right\}.$ (94)
@@
Finally, the dependence of the heat diffusive coefficient $\mu^{*}$ on $\alpha$ is shown in Fig. 5 for three different densities. Note that the coefficient $\mu^{*}$ vanishes in the low-density regime ($\phi=0$) when one neglects the contribution of the kurtosis $a_{2,\text{s}}$ since $\mu^{*}\propto a_{2,\text{s}}$ in this regime BBMG15 . Thus, in order to compare the kinetic contribution to $\mu^{*}$ when $\phi=0$ with the collisional contributions to this coefficient for dense gases ($\phi\neq 0$), the three curves considered in Fig. 5 have been obtained from the results displayed in Table 1 but including the pure dilute contribution to $\mu^{*}$ coming from $a_{2,\text{s}}$. This (new) contribution is given by $a_{2,\text{s}}/2\nu_{\kappa}^{*}$ for $d=2$ BBMG15 . Note also that the collision integral (94) appearing in the expression of $\mu^{*}$ has been still computed by assuming $a_{2,\text{s}}=0$ for the sake of simplicity. It is quite apparent that the impact of density on $\mu^{*}$ is quite significant since while this coefficient is always negative for moderately dense gases ($\phi\lesssim 0.3$), it can be positive at higher densities for $\alpha\lesssim\alpha_{0}$ where the value of $\alpha_{0}$ depends on the volume fraction $\phi$. The value of $\alpha_{0}$ increases with density: for instance, $\alpha_{0}\simeq 0.33$ for $\phi=0.4$. Moreover, as expected, the magnitude of $\mu^{*}$ is quite small for any density. This means that, for practical purposes, one can neglect the contribution to the heat flux coming from the term proportional to the density gradient, namely, the heat flux obeys Fourier’s law $\mathbf{q}^{(1)}=-\kappa\nabla T$. This conclusion contrasts with the results obtained for undriven granular fluids GD99 where the (reduced) heat diffusive coefficient $\mu^{*}$ is clearly different from zero for strong collisional dissipation.
@@ Acknowledgements.
×$\times\left[\psi(\mathbf{v}_{1}^{\prime})-\psi(\mathbf{v}_{1})\right]f_{2}(\mathbf{r},\mathbf{v}_{1},\mathbf{r}+\boldsymbol{\sigma},\mathbf{v}_{2},t),$
@@
×$\times\left[\psi(\mathbf{v}_{2}^{\prime})-\psi(\mathbf{v}_{2})\right]f_{2}(\mathbf{r},\mathbf{v}_{2},\mathbf{r}-\boldsymbol{\sigma},\mathbf{v}_{1},t).$
@@
×$\times\int_{0}^{1}\;\text{d}\lambda\;f_{2}\left[\mathbf{r}-\lambda\boldsymbol{\sigma},\mathbf{v}_{1},\mathbf{r}+(1-\lambda)\boldsymbol{\sigma},\mathbf{v}_{2},t)\right].$
@@
[…]
@@
×$\times\int_{0}^{1}\;\text{d}\lambda\;f_{2}\left[\mathbf{r}-\lambda\boldsymbol{\sigma},\mathbf{v}_{1},\mathbf{r}+(1-\lambda)\boldsymbol{\sigma},\mathbf{v}_{2},t)\right],$
@@
×$\times\int_{0}^{1}\;\text{d}\lambda\;f_{2}\left[\mathbf{r}-\lambda\boldsymbol{\sigma},\mathbf{v}_{1},\mathbf{r}+(1-\lambda)\boldsymbol{\sigma},\mathbf{v}_{2},t)\right],$

### v7

@@ I Introduction
$\kappa_{\text{k}}^{*}=\frac{1}{2\nu_{\kappa}^{*}+\Delta_{s}^{*}\Big(\frac{\partial\zeta_{0}^{*}}{\partial\Delta^{*}}\Big)_{s}}\left\{1+\frac{3}{8}\phi\chi(1+\alpha)^{2}(2\alpha-1)-\frac{\Delta_{\text{s}}^{*}}{\sqrt{2\pi}}\phi\chi\left[\frac{3}{4}+3(1+\alpha)\left(1-\sqrt{\frac{\pi}{2}}\Delta_{\text{s}}^{*}\right)-\frac{9}{2}(1+\alpha)^{2}-\Delta_{\text{s}}^{*2}\right]\right\}$,
@@
$\mu_{\text{k}}^{*}=-\frac{1}{\nu_{\kappa}^{*}}\phi\chi\left(1+\frac{1}{2}\phi\partial_{\phi}\ln\chi\right)\left\{\frac{3}{8}\alpha(1-\alpha^{2})-\frac{\Delta_{\text{s}}^{*}}{\sqrt{2\pi}}\phi\chi\left[2\Delta_{\text{s}}^{*2}-3\left(\frac{1}{2}-\alpha^{2}\right)+\frac{3}{2}\sqrt{2\pi}\alpha\Delta_{\text{s}}^{*}\right]\right\}$,
@@ II.2 Enskog kinetic equation
$\times(-\widehat{\boldsymbol{\sigma}}\cdot{\bf g}-2\Delta)\alpha^{-2}\chi(\mathbf{r},\mathbf{r}+\boldsymbol{\sigma})f(\mathbf{r},\mathbf{v}_{1}^{\prime\prime},t)$
$\times f(\mathbf{r}+\boldsymbol{\sigma},\mathbf{v}_{2}^{\prime\prime},t)-\sigma^{d-1}\int\ \text{d}{\bf v}_{2}\int\text{d}\widehat{\boldsymbol{\sigma}}\Theta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})$
$\times(\widehat{\boldsymbol{\sigma}}\cdot{\bf g})\chi(\mathbf{r},\mathbf{r}+\boldsymbol{\sigma})f(\mathbf{r},\mathbf{v}_{1},t)f(\mathbf{r}+\boldsymbol{\sigma},\mathbf{v}_{2},t).$ (9)
@@
$\times\chi(\mathbf{r},\mathbf{r}+\boldsymbol{\sigma})f(\mathbf{r},\mathbf{v}_{1},t)f(\mathbf{r}+\boldsymbol{\sigma},\mathbf{v}_{2},t)\left[\psi(\mathbf{v}_{1}^{\prime})-\psi(\mathbf{v}_{1})\right],$
@@ III.1 Zeroth-order approximation
$\times(-\widehat{\boldsymbol{\sigma}}\cdot{\bf g}-2\Delta)\alpha^{-2}f^{(0)}(\mathbf{v}_{1}^{\prime\prime})f^{(0)}(\mathbf{v}_{2}^{\prime\prime})-\sigma^{d-1}\chi$
$\times\int\ \text{d}{\bf v}_{2}\int\text{d}\widehat{\boldsymbol{\sigma}}\Theta(\widehat{{\boldsymbol{\sigma}}}\cdot{\bf g})(\widehat{\boldsymbol{\sigma}}\cdot{\bf g})f^{(0)}(\mathbf{v}_{1})f^{(0)}(\mathbf{v}_{2}).$ (29)
@@
$\zeta^{(0)}T\partial_{T}f^{(0)}=J_{\text{E}}^{(0)}[f^{(0)},f^{(0)}].$ (32)
@@
$\times\left(B_{1}g^{*}\Delta^{*2}+B_{2}\alpha g^{*2}\Delta^{*}-\frac{1-\alpha^{2}}{4}B_{3}g^{*3}\right),$
@@ III.2 First-order approximation
[…]
@@
${\bf A}\left(\mathbf{V}\right)$ $=$ $-\mathbf{V}T\frac{\partial f^{(0)}}{\partial T}-\frac{p}{\rho}\left(1+T\frac{\partial}{\partial T}\ln p^{*}\right)\frac{\partial f^{(0)}}{\partial\mathbf{V}}$ (47)
$-\boldsymbol{\mathcal{K}}\left[T\frac{\partial f^{(0)}}{\partial T}\right],$
${\bf B}\left(\mathbf{V}\right)$ $=$ $-{\bf V}f^{(0)}-\frac{p}{\rho}\left(1+\phi\frac{\partial}{\partial\phi}\ln p^{*}\right)\frac{\partial f^{(0)}}{\partial\mathbf{V}}$ (48)
$-\left(1+\frac{1}{2}\phi\frac{\partial}{\partial\phi}\ln\chi\right)\boldsymbol{\mathcal{K}}\left[f^{(0)}\right],$
@@
$D\left(\mathbf{V}\right)$ $=$ $\frac{1}{d}\frac{\partial}{\partial\mathbf{V}}\cdot\left(\mathbf{V}f^{(0)}\right)+\left(\zeta_{U}+\frac{2}{d}p^{*}\right)T\frac{\partial f^{(0)}}{\partial T}$ (50)
$+\frac{1}{d}\mathcal{K}_{i}\left[\frac{\partial f^{(0)}}{\partial V_{i}}\right].$
@@
$\times(-\widehat{\boldsymbol{\sigma}}\cdot\mathbf{g}-2\Delta)\widehat{\sigma}_{i}\alpha^{-2}f^{(0)}(\mathbf{v}_{1}^{\prime\prime})X(\mathbf{v}_{2}^{\prime\prime})$
@@
$\times f^{(0)}(\mathbf{v}_{1})X(\mathbf{v}_{2}).$
@@ IV.1 Pressure tensor
$I_{\eta}$ $=\int\text{d}\mathbf{c}_{1}\int\text{d}\mathbf{c}_{2}\;g^{*-1}g_{x}^{*2}g_{y}^{*2}\varphi_{\text{M}}(\mathbf{c}_{1})\varphi_{\text{M}}(\mathbf{c}_{2}),$ (57)
@@ IV.2 Heat flux
$\times\frac{1}{dn}\int\;\text{d}\mathbf{v}\mathbf{S}(\mathbf{V})\cdot\boldsymbol{\mathcal{K}}\left[f^{(0)}\right],$
@@
[…]
@@
[…]
@@
$I_{\eta}=\frac{3}{8}\sqrt{\frac{\pi}{2}},\quad I_{\gamma}=\sqrt{\frac{\pi}{2}},\quad{I}_{\eta}^{\prime}=-\sqrt{\frac{\pi}{2}}\left(1+2\alpha\right),$ (88)
@@
[…]
@@
$\int\text{d}{\bf V}{\bf S}({\bf V})\cdot\boldsymbol{\mathcal{K}}\left[\frac{\partial}{\partial\mathbf{V}}\cdot\left(\mathbf{V}f^{(0)}\right)\right]=$ $8\phi\chi\frac{nT^{2}}{m}\bigg\{\frac{3}{8}(1+\alpha)^{2}(1-2\alpha)+\frac{\Delta^{*}}{\sqrt{2\pi}}\bigg[\frac{3}{4}+3(1+\alpha)\left(1-\sqrt{\frac{\pi}{2}}\Delta^{*}\right)$
$-\frac{9}{2}(1+\alpha)^{2}-\Delta^{*2}\bigg]\bigg\},$ (93)
$\int\text{d}\mathbf{v}\mathbf{S}(\mathbf{V})\cdot\boldsymbol{\mathcal{K}}\left[f^{(0)}\right]=$ $-8\phi\chi\frac{nT^{2}}{m}\left\{\frac{3}{8}\alpha(1-\alpha^{2})-\frac{\Delta^{*}}{\sqrt{2\pi}}\left[2\Delta^{*2}-3\left(\frac{1}{2}-\alpha^{2}\right)+\frac{3}{2}\sqrt{2\pi}\alpha\Delta^{*}\right]\right\}.$ (94)
@@
Finally, the dependence of the heat diffusive coefficient $\mu^{*}$ on $\alpha$ is shown in Fig. 5 for three different densities. Note that the coefficient $\mu^{*}$ vanishes in the low-density regime ($\phi=0$) when one neglects the contribution of the kurtosis $a_{2,\text{s}}$ since $\mu^{*}\propto a_{2,\text{s}}$ in this regime BBMG15 . Thus, in order to compare the kinetic contribution to $\mu^{*}$ when $\phi=0$ with the collisional contributions to this coefficient for dense gases ($\phi\neq 0$), the three curves considered in Fig. 5 have been obtained from the results displayed in Table 1 but including the pure dilute contribution to $\mu^{*}$ coming from $a_{2,\text{s}}$. This (new) contribution is given by $a_{2,\text{s}}/2\nu_{\kappa}^{*}$ for $d=2$ BBMG15 . Note also that the collision integral (94) appearing in the expression of $\mu^{*}$ has been still computed by assuming $a_{2,\text{s}}=0$ for the sake of simplicity. It is quite apparent that this coefficient is always negative and the impact of density on $\mu^{*}$ is in general quite significant. Moreover, as expected, the magnitude of $\mu^{*}$ is quite small for any density. This means that, for practical purposes, one can neglect the contribution to the heat flux coming from the term proportional to the density gradient, namely, the heat flux obeys Fourier’s law $\mathbf{q}^{(1)}=-\kappa\nabla T$. This conclusion contrasts with the results obtained for undriven granular fluids GD99 where the (reduced) heat diffusive coefficient $\mu^{*}$ is clearly different from zero for strong collisional dissipation.
@@ Acknowledgements.
$\times\left[\psi(\mathbf{v}_{1}^{\prime})-\psi(\mathbf{v}_{1})\right]f_{2}(\mathbf{r},\mathbf{v}_{1},\mathbf{r}+\boldsymbol{\sigma},\mathbf{v}_{2},t),$
@@
$\times\left[\psi(\mathbf{v}_{2}^{\prime})-\psi(\mathbf{v}_{2})\right]f_{2}(\mathbf{r},\mathbf{v}_{2},\mathbf{r}-\boldsymbol{\sigma},\mathbf{v}_{1},t).$
@@
$\times\int_{0}^{1}\;\text{d}\lambda\;f_{2}\left[\mathbf{r}-\lambda\boldsymbol{\sigma},\mathbf{v}_{1},\mathbf{r}+(1-\lambda)\boldsymbol{\sigma},\mathbf{v}_{2},t)\right].$
@@
$I_{\frac{m}{2}v^{2}}$ $=$ $\frac{m}{2}\sigma^{d-1}\int\text{d}\mathbf{v}_{1}\int\text{d}\mathbf{v}_{2}\int d\widehat{\boldsymbol{\sigma}}\,\Theta(\widehat{{\boldsymbol{\sigma}}}\cdot\mathbf{g})(\widehat{\boldsymbol{\sigma}}\cdot{\bf g})\left[\Delta^{2}+\alpha\Delta(\widehat{\boldsymbol{\sigma}}\cdot{\bf g})-\frac{1-\alpha^{2}}{4}(\widehat{\boldsymbol{\sigma}}\cdot{\bf g})^{2}\right]f_{2}(\mathbf{r},\mathbf{v}_{1},\mathbf{r}+\boldsymbol{\sigma},\mathbf{v}_{2},t)$ (106)
@@
$\times\int_{0}^{1}\;\text{d}\lambda\;f_{2}\left[\mathbf{r}-\lambda\boldsymbol{\sigma},\mathbf{v}_{1},\mathbf{r}+(1-\lambda)\boldsymbol{\sigma},\mathbf{v}_{2},t)\right],$
@@
$\times\int_{0}^{1}\;\text{d}\lambda\;f_{2}\left[\mathbf{r}-\lambda\boldsymbol{\sigma},\mathbf{v}_{1},\mathbf{r}+(1-\lambda)\boldsymbol{\sigma},\mathbf{v}_{2},t)\right],$